solve the following problem, be sure to show the equation used for the solution.
The lenght of a rectangle is 2in. more than twice its width. if the perimeter of the retangle is 34 in., find the dimensions of the rectangle.

Answer by venugopalramana(3286)

solve the following problem, be sure to show the equation used for the solution.
The lenght....L SAY
of a rectangle is 2in. more than twice its width....W SAY
TWICE WIDTH =2W
2" MORE=2W+2
L=2W+2
PERIMETER=2(L+W)=2(2W+2+W)=34
3W+2=34/2=17
3W=17-2=15
W=15/3=5
L=2W+2=2*5+2=12
if the perimeter of the retangle is 34 in., find the dimensions of the rectangle.